# Leveraging AI-Generated Emotional Self-Voice to Nudge People towards their Ideal Selves

### v1

@@ Abstract.
Emotions, shaped by past experiences, significantly influence decision-making and goal pursuit. Traditional cognitive-behavioral techniques for personal development rely on mental imagery to envision ideal selves, but may be less effective for individuals who struggle with visualization. This paper introduces Emotional Self-Voice (ESV), a novel system combining emotionally expressive language models and voice cloning technologies to render customized responses in the user’s own voice. We investigate the potential of ESV to nudge individuals towards their ideal selves in a study with 60 participants. Across all three conditions (ESV, text-only, and mental imagination), we observed an increase in resilience, confidence, motivation, and goal commitment, but the ESV condition was perceived as uniquely engaging and personalized. We discuss the implications of designing generated self-voice systems as a personalized behavioral intervention for different scenarios.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information systems applicationsccs: Human-centered computing Natural language interfaces
Figure 1. The Emotional Self Voice (ESV) system renders emotionally expressive ideal-self responses to difficult situations in the user’s own voice using emotionally expressive language models and voice cloning technologies. Left: in the example of an individual facing the failure to achieve a goal, he hears his ideal self speaking to himself in his voice, acknowledging the setbacks and reframing the failure as an opportunity to learn. Right: in another example of the individual struggling to keep up with a personal habit, she hears her ideal self speaking with her voice, embodying the habit as her identity as a form of motivation.
@@ 1. Introduction
Emotion plays a crucial role in shaping our decisions and behaviors, acting as a manifestation of our predictions based on prior experiences (Barrett, 2017). As a result, emotion affects how we make decisions. For example, a positive outlook on the future can foster resilience and grit, enabling us to persevere in the face of challenges, but a pessimistic perspective may deter us from pursuing our goals.
Within this context, the concept of an ideal or aspirational version of one’s future self, rooted in self-discrepancy theory, has been shown to be a key driver of intrinsic motivation. It provides a powerful framework for understanding goal-directed behavior and guides regulatory systems to reduce the perceived gap between one’s current or actual self and the ideal self by making decisions consistent with one’s goals (Higgins et al., 1994). In the same vein, research has shown that the vividness of the ”future self” can promote future-oriented behaviors and reduce the gap between intention and behavior (Hershfield et al., 2011; Rösch et al., 2022). Traditional cognitive-behavioral therapy (CBT) and reframing practices aim at helping individuals set goals and receive positive reinforcement for their efforts often encourage individuals to create mental imagery of their alternative versions of themselves in a scenario. However, producing compelling mental images (including images, voice and other sensory modalities) can be difficult, and some lack the ability altogether (Andrade et al., 2014). The question is then how to nudge an individual towards their ideal self. Concretely, the challenges lie in one’s ability to experience an ideal version of themselves that is temporally and emotionally distant from their current self.
@@
Voice presents a particularly intriguing avenue for exploration: common phrases like ”inner voice” and ”voices in one’s head” reflect the natural tendency to mentally simulate thoughts as speech, and hearing one’s voice elicits higher arousal and alertness in a way that may be beneficial for goal achievement (Payne et al., 2021; Kim and Song, 2024). The advent of large language model (LLM) enabled chatbots that can converse with users in natural language through both text and speech also opens up compelling possibilities for their positive application in behavioral interventions.
@@
Our research aims to investigate how hearing one’s own voice respond to challenging situations with different attitudes affects behavior, confidence, and motivation for goal achievement. We hypothesize that: (1) Hearing an ideal-self response through ESV is more motivating and confidence-boosting than text-only or control conditions, despite potential initial discomfort with hearing one’s own voice; (2) ESV enables participants to more vividly imagine their ideal-self saying the response; (3) ESV leads to greater improvements in affect, resilience, and goal commitment compared to other conditions.
To investigate the potential of ESV, we conducted a between-subjects study with 60 participants with random assignment to one of three conditions: ESV, a text-only variant condition (seeing the generated ideal-self response in text without hearing the generated self-voice), or control (only mental imagination of the ideal-self response). The study focused on two scenarios related to goal achievement: facing failure in achieving a goal and experiencing difficulty in establishing a habit.
@@
    The ESV system successfully generated highly realistic, emotionally expressive self-voices, as validated by our technical characterization and participant feedback.
@@
    Across all conditions, we observed an increase in emotional valence, resilience, confidence, motivation, and commitment. However, the ESV condition showed unique benefits in certain areas.
@@
    The ideal self concept, as presented through ESV, demonstrated potential in enhancing the vividness of future-oriented thinking and goal representation.
@@
  5. (5)
    The study revealed nuanced differences between past-focused (failure) and future-focused (habit formation) scenarios, highlighting the importance of temporal framing in behavioral interventions.
@@
  • A novel pipeline for generating Emotional Self-Voice using generative text-to-speech and speech-to-speech models.
  • Technical validation of the emotional and speech quality of the generated self-voice.
  • Comprehensive user study results that provide insights into the perception and effectiveness of self-voice interventions.
  • Implications for self-discrepancy theory and the role of ideal self representations in behavioral change.
  • Design implications for future self-voice systems as a form of personalized behavioral intervention.
@@ 2.1. Towards the Ideal Self
The ideal self is defined as “a personal vision, or an image of what kind of person one wishes to be” (Boyatzis and Akrivou, 2006). A breadth of positive psychology research suggests that the ideal-self concept serves as a key driver of intrinsic motivation to achieve personal and professional goals (Higgins, 2024; Yeager and Dweck, 2020), by providing a positive reference value that guides regulatory systems to eventually align one’s current or actual self with the ideal, aspirational self (Higgins et al., 1994). Developing ideal-self concepts also implicitly suggests that the current self is malleable and that qualities of the desired self can be developed, a mindset that promotes resilience and achievement (Yeager and Dweck, 2012). Apart from positive references, self-regulation can also occur through avoidance of undesired end states: for example, commitment devices such as financial penalties and social accountability have shown a positive relationship with goal achievement (Lee et al., 2021). However, prior work has shown that the ideal self tends to show a stronger positive association with desired end states than a negative association with undesired end states, suggesting that movement toward the ideal self benefits from approach systems rather than avoidance systems (Higgins et al., 1994).
Existing research on interventions for behavior change and goal achievement have primarily explored their potential to support goal capture, goal monitoring and maintenance of goal-directed motivation (Lolla and Sas, 2023). Despite the popularity of such interventions, with over 1,300 apps related to goal achievement available on the Google Play Store, they often suffer from high attrition and limited user engagement (Lazar et al., 2015). To address these issues, Dominick and Cole (2020) proposed the incorporation of goals into one’s identity and found that framing a goal as part of one’s identity (e.g., “I will think of myself as someone who eats healthy and works out”) promotes goal-consistent choices to a greater extent than simply setting the goal (e.g., “I will be mindful about what I eat”). Reflection has also emerged as an important component of behavior change interventions (Cho et al., 2022), with studies finding that it not only facilitates goal monitoring, but also helps users see positive aspects of initially negative experiences, and draw lessons for future behavior (Isaacs et al., 2013).
@@ 2.2. AI-generated Synthetic Selves for Behavior Change
Digital representations of oneself can alter our behavior. The Proteus effect (Yee and Bailenson, 2007) is a phenomenon in which people’s behavior adapt to the traits of their avatars in virtual worlds. In virtual reality (VR), it has been shown that embodying different characters can influence people’s cognition and behavior. For example, men who embodied Einstein in VR performed better on a cognitive task (Banakou et al., 2018), and embodying a young child-like avatar in VR can lead to changes in perception of object size and speaking behaviors (Tajadura-Jiménez et al., 2017). In a study done by Osimo et al., subjects embodied Sigmund Freud and offered themselves counseling, capitalizing on the illusion of body ownership of the Freud body (Osimo et al., 2015).
Beyond static avatars, recent advancements in machine learning and generative AI technologies have made it possible to produce synthetic media for the purposes of wellbeing, learning, and skill development (Pataranutaporn et al., 2021; Leong, 2023; Danry et al., 2022). Researchers for instance have explored improving engagement in exercise (Clarke et al., 2023) or developing public speaking skills by generating personalized videos of oneself speaking confidently in public speaking scenarios (Clarke et al., 2023; Leong, 2021). Synthesized future representations of oneself have also been shown to improve people’s tendency to save for retirement (Hershfield et al., 2011), and chatting with an AI-generated future version of oneself can positively influence emotions and a feeling of self-continuity (Pataranutaporn et al., 2024). In contrast to generated videos, augmented reality (AR) filters have been shown as a means for real-time interventions. For example, AR filters making oneself look more inventor- or child-like has shown to possibly boost creativity (Leong et al., 2021). In another example, it was found that the private application of filters on oneself in online public speaking situations can be a means to mitigate public speaking anxiety and encourage further development of the skill (Leong et al., 2023).
More related to our work are the explorations into synthetic voices. Costa et al. (2018) found that by playing back an altered version of one’s voice to be calmer made people feel less anxious during relationship conflicts, whereas lowering the voice made people feel more powerful during debates. Explorations have also been made into using the voice of oneself (Kim and Song, 2024) or familiar people (e.g., family and friends) to improve the efficacy of notifications (Chan et al., 2021).
Next, we discuss prior work that use voice as a behavioral intervention and the generation of emotionally expressive voices.
@@ 2.3. Voice-based Intervention for Emotion Regulation
Hearing the playback of one’s own voice is not common, and most people find it unpleasant (Holmes et al., 2018; Kim and Sundar, 2012) and it could cause attentional bias (Daryadar and Raghibi, 2015). People also had negative affective and defensive reactions to hearing one’s voice due to the mismatch in what they expected to hear and what was heard (Holzman and Rousey, 1966; Holzman et al., 1966). Voices not only express feelings but can also affect one’s feelings (Aucouturier et al., 2016; Hatfield and Hsee, 1995). Using one’s own voice as an intervention has a few favorable aspects. Perceptually, hearing one’s voice elicits higher arousal and alertness (Payne et al., 2021), and are possibly more familiar than other voices (Xu et al., 2013). Self-voices also seem to be processed implicitly, even when they are unrecognized by their owners (Daryadar and Raghibi, 2015). Psychologically, there are a few effects that work in the favor of using self-voice. First, people tend to process self-related information more importantly and positively due to the self-referencing effect (e.g., when the information states things about or related to “you”) (Rogers et al., 1977). Self-referencing message have been shown to improve involvement and intention to increase physical activity and healthy eating (Wirtz, 2020). Many voices seem familiar due to frequent exposure or the “mere exposure” effect that influences personal preferences and enhances affect-based trust (Kwan et al., 2015). Familiarity of voices induces more attention, engagement, and recall (Holmes et al., 2018; Chan et al., 2021; England, 2021). Finally, people might prefer to interact with others (human and digital systems) who are similar to themselves due to the similar-attraction effect (Reeves and Nass, 1996; Nass et al., 1995). These might help strengthen one’s acceptance and embodiment of the emotional self-voice of the ideal self.
@@
In HCI and education research, people reacted most quickly to self-voice notifications compared familiar voices (of instructors and lab mates) and unfamiliar voices (Bhatia and McCrickard, 2006). Self-voice alarms from smartphones produced discomfort but better vocabulary learning task completion and goal achievement compared to other voice alarms (Kim and Song, 2024). Echoing one’s self-voice during learning improves recall (Lane and Chinn, 1986) and developing a second language self-voice improves language learning (Tomlinson, 2001).
@@
However, most of the prior work used individuals’ recorded voices. Only a few studies have explored the digital altering self-voices to influence emotions. A key example was a study that generated a calmer version of one’s voice, which lowered anxiety (Costa et al., 2018). The effects of synthesized versus recorded voices’ effect on emotion and perceptions on content have been studied (Nass et al., 2001), showing that happy synthesized and recorded voices made content seem happier while sad ones made content seem less happy. People also liked the content more and found the information more credible when the emotion of the voice and the content were matched.
@@
More recently, the use of generative AI models for creating emotionally expressive voices has surged, and the models have become increasingly more capable at generating voices with a wide emotional range. Models like EmoCtrl-TTS (Wu et al., 2024) and VAW-GAN (Zhou et al., 2021) can produce time-varying emotional states within the generated speech (Triantafyllopoulos et al., 2023). There is a growing number of open-sourced (Eren and The Coqui TTS Team, 2021; Jia et al., 2018) and proprietary models [^1] https://themetavoice.xyz/  [^2] https://speech.fish.audio/ that enable zero-shot voice cloning.
@@
Researchers have not thoroughly studied the effects of AI generated self voices on emotions and behaviours. Leveraging these emerging models’ capabilities, we can allow the speech (both the content and the voice) to be dynamically generated instead of pre-recorded voices, and we can generate synthetic voices with emotional expressiveness of natural voices. This is an extension of “voice as a design material” (Sutton et al., 2019) where we explore not only the voice characteristics but the combined effect of emotional voice and self voice as an intervention.
@@
In our work, we leverage the latest models for synthesizing both the content from the perspective of an ideal self as well as the spoken voices in the style of the user’s voice. To the best of our knowledge, this work is the first to ever explore the development of and effect using of synthetic emotional, self-voice on individuals’ perception and behavior.

### v2

@@ Abstract.
Emotions, shaped by past experiences, significantly influence decision-making and goal pursuit. Traditional cognitive-behavioral techniques for personal development rely on mental imagery to envision ideal selves, but may be less effective for individuals who struggle with visualization. This paper introduces Emotional Self-Voice (ESV), a novel system combining emotionally expressive language models and voice cloning technologies to render customized responses in the user’s own voice. We investigate the potential of ESV to nudge individuals towards their ideal selves in a study with 60 participants. Across all three conditions (ESV, text-only, and mental imagination), we observed an increase in resilience, confidence, motivation, and goal commitment, and the ESV condition was perceived as uniquely engaging and personalized. We discuss the implications of designing generated self-voice systems as a personalized behavioral intervention for different scenarios.
@@
journalyear: 2025conference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713359isbn: 979-8-4007-1394-1/25/04ccs: Information systems Information systems applicationsccs: Human-centered computing Natural language interfaces
Figure 1. The Emotional Self Voice (ESV) system renders emotionally expressive ideal-self responses to difficult situations in the user’s own voice using emotionally expressive language models and voice cloning technologies. Left: The individual hears his ideal self reframing failure as a learning opportunity. Right: Her ideal self reinforces a habit as part of her identity for motivation.
@@ 1. Introduction
Emotion acts as a manifestation of our predictions based on prior experiences (Barrett, 2017). As a result, emotion affects how we make decisions. For example, a positive outlook on the future can foster resilience and grit, enabling us to persevere in the face of challenges, but a pessimistic perspective may deter us from pursuing our goals. Relatedly, he concept of an ideal or aspirational version of one’s future self, rooted in self-discrepancy theory, has been shown to be a key driver of intrinsic motivation (Higgins et al., 1994). It provides a powerful framework for understanding goal-directed behavior to reduce the perceived gap between one’s current or actual self and the ideal self by making decisions consistent with one’s goals (Higgins et al., 1994). In the same vein, research has shown that the vividness of the “future self” can promote future-oriented behaviors and reduce the gap between intention and behavior (Hershfield et al., 2011; Rösch et al., 2022). Traditional cognitive-behavioral therapy (CBT) and reframing practices aim at helping individuals set goals and receive positive reinforcement for their efforts often encourage individuals to create mental imagery of their alternative versions of themselves in a scenario. However, producing compelling mental images (including images, voice, and other sensory modalities) can be difficult, and some lack the ability altogether (Andrade et al., 2014). The question is then how to nudge an individual towards their ideal self. Concretely, the challenges lie in one’s ability to experience an ideal version of oneself that is temporally and emotionally distant from their current self.
@@
Voice presents a particularly intriguing avenue for exploration: common phrases like ”inner voice” and ”voices in one’s head” reflect people’s natural tendency to mentally simulate thoughts as speech, and hearing one’s voice elicits higher arousal and alertness in a way that may be beneficial for goal achievement (Payne et al., 2021; Kim and Song, 2024). The advent of large language model (LLM) enabled chatbots that can converse with users in natural language through both text and speech also opens up compelling possibilities for their positive application in behavioral interventions.
@@
Our research aims to investigate how hearing one’s own voice respond to challenging situations with different attitudes affects self-perception and behavior for goal achievement. We hypothesize that: (1) Hearing an ideal-self response through ESV is more motivating and confidence-boosting than text-only or mental-imagination-only conditions, despite the potential initial discomfort with hearing one’s own voice; (2) ESV enables people to more vividly imagine their ideal-self saying the response; (3) ESV leads to greater improvements in affect, resilience, and goal commitment compared to other conditions.
To investigate the potential of ESV, we conducted a between-subjects study with 60 participants with random assignment to one of three conditions: ESV, a text-only variant condition (seeing the generated ideal-self response in text without hearing the generated self-voice), or control (only mental imagination of the ideal-self response). The study focused on two scenarios related to goal achievement: overcoming failure in achieving a past goal and experiencing resistance in establishing a habit.
@@
    Overall, the emotional valence ratings became significantly more positive (p¡.001) across most conditions for both scenarios, with the ESV condition having the most positive post-intervention valence ratings. Across all conditions, we observed an increase in resilience, confidence, motivation, and commitment, with the ESV condition again ranked the highest for post-intervention confidence and motivation.
@@
    Qualitative results show that ESV condition has the most positive change in affective states (indicated by the sentiment analysis of post-intervention responses). The ideal-self concept, as presented through ESV, demonstrated the potential to enhance the vividness of future-oriented thinking and goal representation.
@@
    The study revealed nuanced differences between responses towards past-focused unmet outcomes (scenario 1) and future-focused, progress-driven (scenario 2), highlighting the need for different temporal and motivational framing in behavioral interventions.
@@
  • We developed a novel proof-of-concept system that generates Emotional Self-Voice—emotionally expressive responses spoken in the user’s own voice—to help people more vividly embody their ideal self in response to challenging situations.
  • Our system combines generative text-to-speech and speech-to-speech models to successfully generate highly realistic, emotionally expressive self-voices, as validated by our technical characterization and participant feedback.
  • Our user study delved into two scenarios of goal achievement and investigated how ESV affects people’s attitudes towards these personal experiences.
  • Our approach is based on the self-discrepancy theory and we investigate how the concept of an ideal self can be used as an intervention for behavior change. We showed personalized emotional voices can serve as a novel mechanism for bridging the gap between actual and ideal self-perceptions.
  • We also discuss design implications that can be generalized to future self-voice systems as a form of personalized behavioral intervention.
@@ 2.1. Towards the Ideal Self
The ideal self is defined as “a personal vision, or an image of what kind of person one wishes to be” (Boyatzis and Akrivou, 2006). A breadth of positive psychology research suggests that the ideal-self concept serves as a key driver of intrinsic motivation to achieve personal and professional goals (Higgins, 2024; Yeager and Dweck, 2020), by providing a positive reference value that guides regulatory systems to eventually align one’s current or actual self with the ideal, aspirational self (Higgins et al., 1994). Developing ideal-self concepts implicitly suggests that the current self is malleable and that qualities of the desired self can be developed. This perspective closely aligns with research on growth mindsets (Dweck and Yeager, 2019), which finds that the belief that one’s abilities are not fixed but can be improved over time promotes resilience and achievement (Yeager and Dweck, 2012). Apart from positive references, self-regulation can also occur through avoidance of undesired end states: for example, commitment devices such as financial penalties and social accountability have shown a positive relationship with goal achievement (Lee et al., 2021). However, prior work has shown that the ideal self tends to show a stronger positive association with desired end states than a negative association with undesired end states, suggesting that movement toward the ideal self benefits from approach-oriented strategies than avoidance-based ones (Higgins et al., 1994).
Existing research on interventions for behavior change and goal achievement has primarily explored their potential to support goal capture, goal monitoring, and maintenance of goal-directed motivation (Lolla and Sas, 2023; Duckworth et al., 2007). Despite the popularity of such interventions, with over 1,300 apps related to goal achievement available on the Google Play Store, they often suffer from high attrition and limited user engagement (Lazar et al., 2015). To address these issues, Dominick and Cole (2020) proposed the incorporation of goals into one’s identity and found that framing a goal as part of one’s identity (e.g., “I will think of myself as someone who eats healthy and works out.”) promotes goal-consistent choices to a greater extent than simply setting the goal (e.g., “I will be mindful about what I eat.”). Reflection has also emerged as an important component of behavior change interventions (Cho et al., 2022). Applications for technology-mediated reflection, such as reflective conversational agents (Li et al., 2023), tools for recording and reflecting on daily activities (Isaacs et al., 2013) and LLM-assisted journaling (Nepal et al., 2024; Song et al., 2024; Kim et al., 2024) not only facilitate goal monitoring, but also helps users see positive aspects of initially negative experiences, and draw lessons for future behavior.
@@ 2.2. AI-generated Synthetic Selves for Behavior Change
Digital representations of oneself can alter our behavior. The Proteus effect (Yee and Bailenson, 2007) is a phenomenon in which people’s behaviors adapt to the traits of their avatars in virtual worlds. In virtual reality (VR), it has been shown that embodying different characters can influence people’s cognition and behavior. For example, men who embodied Einstein in VR performed better on a cognitive task (Banakou et al., 2018), and embodying a young child-like avatar in VR can lead to changes in perception of object size and speaking behaviors (Tajadura-Jiménez et al., 2017). In a study done by Osimo et al., subjects embodied Sigmund Freud and offered themselves counseling, capitalizing on the illusion of body ownership of the Freud body (Osimo et al., 2015).
Beyond static avatars, recent advancements in machine learning and generative AI technologies have made it possible to produce synthetic self-similar media for the purposes of wellbeing, learning, and skill development (Pataranutaporn et al., 2021; Leong, 2023; Danry et al., 2022). Researchers, for instance, have explored improving engagement in exercise (Clarke et al., 2023) or developing public speaking skills by generating personalized videos of oneself speaking confidently in public speaking scenarios (Clarke et al., 2023; Leong, 2021). In the context of human-AI collaboration, studies find that virtual characters with self-similar appearance are perceived as more intelligent and elicited a stronger sense of co-presence, while those with self-similar voice are perceived as more likable and believable (Guo et al., 2024). Synthesized future representations of oneself have also been shown to improve people’s tendency to save for retirement (Hershfield et al., 2011), and chatting with an AI-generated future version of oneself can positively influence emotions and a feeling of self-continuity (Pataranutaporn et al., 2024). In contrast to generated videos, augmented reality (AR) filters have been shown as a means for real-time interventions. For example, AR filters making oneself look more inventor- or child-like has been shown to possibly boost creativity (Leong et al., 2021). In another example, it was found that the private application of filters on oneself in online public speaking situations can be a means to mitigate public speaking anxiety and encourage further development of the skill (Leong et al., 2023).
More related to our work are the explorations into synthetic voices. Costa et al. (2018) found that playing back an altered version of one’s voice to be calmer made people feel less anxious during relationship conflicts, whereas lowering the voice made people feel more powerful during debates. Explorations have also been made into using the voice of oneself (Kim and Song, 2024) or familiar people (e.g., family and friends) to improve the efficacy of notifications (Chan et al., 2021). Next, we discuss prior work that use voice as a behavioral intervention and the generation of emotionally expressive voices.
@@ 2.3. Voice-based Intervention for Emotion Regulation
Voices not only express feelings but can also affect one’s feelings (Aucouturier et al., 2016; Hatfield and Hsee, 1995). Hearing the playback of one’s own voice is not common. Most people find it unpleasant (Holmes et al., 2018; Kim and Sundar, 2012) and it could cause attentional bias (Daryadar and Raghibi, 2015). People also had negative affective and defensive reactions to hearing one’s voice due to the mismatch in what they expected to hear and what was heard (Holzman and Rousey, 1966; Holzman et al., 1966). Using one’s own voice as an intervention has a few favorable aspects. The presence of self-relevant information in a voice, such as self-similarity and self-generation, has been linked to enhanced speed and accuracy of voice perception (Rosi et al., 2024). Perceptually, hearing one’s voice elicits higher arousal and alertness (Payne et al., 2021), and are possibly more familiar than other voices (Xu et al., 2013). Self-voices also seem to be processed implicitly, even when they are unrecognized by their owners (Daryadar and Raghibi, 2015). Psychologically, there are a few effects that work in the favor of using self-voice. People tend to process self-related information more importantly and positively due to the self-referencing effect (e.g., when the information states things about or related to “you”) (Rogers et al., 1977). Self-referencing messages have been shown to improve involvement and intention to increase physical activity and healthy eating (Wirtz, 2020). Many voices seem familiar due to frequent exposure or the “mere exposure” effect that influences personal preferences and enhances affect-based trust (Kwan et al., 2015). Familiarity of voices induces more attention, engagement, and recall (Holmes et al., 2018; Chan et al., 2021; England, 2021). People might prefer to interact with others (human and digital systems) who are similar to themselves due to the similar-attraction effect (Reeves and Nass, 1996; Nass et al., 1995). These might help strengthen one’s acceptance and embodiment of the emotional self-voice of the ideal self.
@@
In HCI and education research, people reacted most quickly to self-voice notifications compared to familiar voices (of instructors and lab mates) and unfamiliar voices (Bhatia and McCrickard, 2006). Video game avatars with self-similar voices elicited greater identification, and improved performance and immersion in the game (Kao et al., 2021). Self-voice alarms from smartphones produced discomfort but better vocabulary learning task completion and goal achievement compared to other voice alarms (Kim and Song, 2024). Echoing one’s self-voice during learning improves recall (Lane and Chinn, 1986) and developing a second language self-voice improves language learning (Tomlinson, 2001).
@@
However, most of the prior work used individuals’ recorded voices. Only a few studies have explored the digital altering of self-voices to influence emotions. A key example was a study that generated a calmer version of one’s voice, which lowered anxiety (Costa et al., 2018). The effects of synthesized versus recorded voices’ effect on emotion and perceptions of content have been studied (Nass et al., 2001), showing that happy synthesized and recorded voices made content seem happier while sad ones made content seem less happy. People also liked the content more and found the information more credible when the emotion of the voice and the content were matched.
@@
More recently, the use of generative AI models for creating emotionally expressive voices has surged, and the models have become increasingly more capable of generating voices with a wide emotional range. Models like EmoCtrl-TTS (Wu et al., 2024) and VAW-GAN (Zhou et al., 2021) can produce time-varying emotional states within the generated speech (Triantafyllopoulos et al., 2023). There is a growing number of open-sourced (Eren and The Coqui TTS Team, 2021; Jia et al., 2018) and proprietary models [^1] https://themetavoice.xyz/  [^2] https://speech.fish.audio/ that enable zero-shot voice cloning.
@@
Researchers have not thoroughly studied the effects of AI generated self voices on emotions and behaviors. Leveraging these emerging models’ capabilities, we can allow the speech (both the content and the voice) to be dynamically generated instead of pre-recorded voices, and we can generate synthetic voices with the emotional expressiveness of natural voices. This is an extension of “voice as a design material” (Sutton et al., 2019) where we explore not only the voice characteristics but the combined effect of emotional voice and self voice as an intervention.
@@
In our work, we leverage the latest models for synthesizing both the content from the perspective of an ideal self as well as the spoken voices in the style of the user’s voice. To the best of our knowledge, this work is the first to ever explore the development and effect of using synthetic emotional, self-voice on individuals’ perception and behavior.